OpenStudy (anonymous):

Which of the following equations is written in slope-intercept form? 2x + 3y = 6 x = -three halvesy + 3 3y = -2x + 6 y = -two thirdsx + 2

OpenStudy (anonymous):

Its the third one: y=-2/3x+2

OpenStudy (britanyashton):

slope intercept form-y=mx+b point slope form- (y-yv1)=m(x-xv1)